A generic salt, AB2 , has a molar mass of 151 g/mol and a solubility of 3.80 g/L at 25 °C.

AB2(s)↽−−⇀A2+(aq)+2B−(aq)

What is the
1
Expert's answer
2020-04-27T01:40:47-0400

Correct task : A generic salt AB2, has a molar mass of 151 g/mol and a solubility of 3.80 g/L at 25oC. What is the Ksp of this salt at 25oC.


Solution:

When the salt dissolves, it dissociates as follows:

AB2 --> A2+ + 2B⁻

--S--------S-------2S--


The molar solubility (S) is the number of moles that can dissolve in 1 L of solution.

Molar solubility of salt is (3.80 g/L) / (151 g/mol) = 0.0252 mol/L.

According to the dissociation equation:

Solubility of A2+ is 0.0252 mol/L and solubility of B⁻ is 2×0.0252 mol= 0.0504 mol/L.

[A2+] = 0.0252 mol/L.

[B⁻] = 0.0504 mol/L.


Ksp is the solubility product constant and calculated as follows:

Ksp(AB2) = [A2+] × [B⁻]2

Ksp(AB2) = [0.0252] × [0.0504]2

Ksp(AB2) = 6.40×10-5.

Ksp of this salt is 6.40×10-5.


Answer: 6.40×10-5 is the Ksp of this salt at 25oC.